# Hybrid-Task Meta-Learning: A Graph Neural Network Approach for Scalable and Transferable Bandwidth Allocation

### v2

@@ Abstract
In this paper, we develop a deep learning-based bandwidth allocation policy that is: 1) scalable with the number of users and 2) transferable to different communication scenarios, such as non-stationary wireless channels, different quality-of-service (QoS) requirements, and dynamically available resources. To support scalability, the bandwidth allocation policy is represented by a graph neural network (GNN), with which the number of training parameters does not change with the number of users. To enable the generalization of the GNN, we develop a hybrid-task meta-learning (HML) algorithm that trains the initial parameters of the GNN with different communication scenarios during meta-training. Next, during meta-testing, a few samples are used to fine-tune the GNN with unseen communication scenarios. Simulation results demonstrate that our HML approach can improve the initial performance by $8.79\%$, and sampling efficiency by $73\%$, compared with existing benchmarks. After fine-tuning, our near-optimal GNN-based policy can achieve close to the same reward with much lower inference complexity compared to the optimal policy obtained using iterative optimization.
@@ I Introduction
A key 5G application that requires flexible resource allocation solutions is network slicing, where resources from a shared physical infrastructure is partitioned into distinct network slices supporting diverse QoS requirements, such as data rate [21, 20], latency [22, 23], and security [25, 26, 24, 27], in both long and short coding blocklength regimes [30, 28, 29]. To reserve resources for a single slice, the authors of [31] proposed to compute the weights of different slices based on the corresponding QoS requirements and the number of service requests. With this approach, the amount of reserved resources for each slice is stochastic. Meanwhile, since the wireless channels are non-stationary, the reserved resources and the wireless channels in the training stage could be different from the actual required resources in the testing stage [32, 33]. As such, the mismatch between training data samples and testing data samples remains a crucial bottleneck for implementing efficient learning-based policies in practical wireless networks.
Recent works have proposed to reduce the online training time by transfer learning, which involves offline pre-training and online fine-tuning [10]. This method effectively reuses previously well-trained neural network features and significantly improves the sample efficiency. To further improve the online training efficiency for unseen tasks, meta-learning has been proposed [34, 37, 35, 36]. One of the meta-learning algorithms, model-agnostic meta-learning (MAML), has been applied to solve policy mismatch issues caused by varying user requests and non-stationary wireless channels [39, 38, 8, 9]. While these aforementioned works have highlighted the generalization ability of meta-learning for non-stationary wireless resource allocation, no works have addressed the impact of diverse QoS requirements in different communication scenarios.
@@
  • Finally, we develop our generalized hybrid-task meta-learning (HML) algorithm that is transferable to different communication scenarios by using meta-training to train the initial parameters of the GNN. We note that only a few samples are required to fine-tune the parameters of the GNN in meta-testing which validates that our GNN-based policy initialized by HML can be efficiently transferred to previously unseen communication scenarios. Simulation results show that our GNN-based policy achieves near-optimal performance and HML significantly outperforms the three considered benchmarks of MAML, MTL transfer (multi-task learning based transfer learning), and random initialisation.
In our simulations, the gap between the sum reward achieved by the GNN-based policy and that of the optimal bandwidth allocation policy obtained from the iterative optimization algorithm is less than $6\%$. HML also improves the initial performance by up to $8.79\%$ and sample efficiency by up to $73\%$ compared with the MAML benchmark. We also show that the performance gains of HML is even higher when compared to the other two benchmarks.
@@ II Related Works
II-A Deep Learning for Resource Allocation in Wireless Communications
Applying deep learning for resource allocation in wireless networks has been widely studied in the existing literature [15, 16]. In [15], the authors showed that learning-based algorithms could obtain near-optimal solutions, and the computational complexity in inference is low. In [16], the authors proposed a FNN-based unsupervised learning algorithm to optimize the bandwidth allocation policy. More recently, due to the fact that FNN is not scalable to the number of users, GNNs have been applied in wireless networks optimizations [19, 18]. In [18], the authors designed a GNN, which is scalable to the number of users in a wireless network, to minimize the summation of queuing delay violation probability and packet loss probability. In [19], the authors developed GNN-based scalable learning-based methods to solve radio resource management problems.
@@ II-A Deep Learning for Resource Allocation in Wireless Communications
Refs QoS Data rate Latency Security
Long Short Long Short Long Short
@@ II-B Generalization of Deep Learning Policies in Non-Stationary Wireless Networks
In wireless networks, the user requests, wireless channels, and available resources for each type of service can be non-stationary. Table I summarizes some QoS requirements considered in the related works. For example, data rate, latency, and security have been investigated in [20, 21, 22, 23, 24, 25, 26]. These papers mainly focus on scenarios with long channel coding blocklengths, where the achievable rate of a wireless link can be approximated by the Shannon capacity. In 5G, the coding blocklength can be short, and Shannon capacity is not applicable. As such, the authors of [27, 28] established how to optimize wireless communication systems using the achievable rate in the short blocklength regime [29]. Meanwhile, different services may co-exist in one network, and the authors of [30, 10] considered different QoS requirements in both long and short blocklength regimes. To support diverse QoS requirements in network slicing, the authors of [31] proposed to reserve bandwidth for different slices based on the number of users and the required QoS.
Further considering that the number of requests, the reserved resources, and the wireless channels are dynamic, improving the generalization ability of deep learning policies has attracted significant research interests in recent years. One approach to address this challenge is to carefully initialize the neural network and fine-tune it online. The authors of [10] applied transfer learning to fine-tune the parameters of deep neural networks that are trained offline in dynamic wireless networks. To further improve the sample efficiency in an unseen communication scenario, meta-learning has been adopted in [8, 9, 38, 39], where the hyper-parameters of a deep neural network, such as the initial parameters, are updated according to a set of communication scenarios in meta-training. In [38], meta-learning was applied to optimize computing resource allocation policies in mobile edge computing networks to fit both time-varying wireless channels and different requests of computing tasks. In [39], meta-learning was applied in virtual reality to quickly adapt to the user movement patterns changing over time. To improve the training efficiency in non-stationary vehicle networks, the authors in [8] proposed optimizing the beamforming using meta-learning. In [9], the authors combined meta-learning and support vector regression to extract the features for beamforming optimization, further improving training efficiency over non-stationary channels.
@@ III System Model and Problem Formulation
III-A Different QoS in Infinite and Short Blocklength Regimes
To investigate the generalization ability of our proposed bandwidth allocation policy, we consider both long and short blocklength regimes with three types of QoS requirements, i.e., data rate, queuing delay, and security. Thus, there are six scenarios in total. We denote the reward of the $u$-th user by
$r_{u}^{\Phi,\xi},\quad\Phi\in\{D,E,S\}\text{ and }\xi\in\{\mathcal{I,F}\},$ (1)
where superscripts $D,E,S$ represent data rate, effective capacity with queuing delay constraint, and secrecy rate, respectively, whilst the superscripts $\mathcal{I,F}$ represent the scenarios in the infinite long and finite short blocklength regimes, respectively.
III-A1 Data Rate Requirement
When the blocklength is long, the data rate reward of the $u$-th user can be expressed as
@@ III-A1 Data Rate Requirement
III-A2 Latency Requirement
When considering latency constraints due to queueing delays, the effective capacity is applied to characterize the statistical QoS requirement in wireless communications, and is expressed as [28]
$\begin{split}r_{u}^{E,\xi}=&-\frac{1}{\vartheta_{u}T_{\mathrm{c}}}\ln\left(\mathbb{E}_{g_{u}}\left[\exp\left(-\vartheta_{u}T_{\mathrm{c}}r_{u}^{D,\xi}\right)\right]\right),\xi\in\{\mathcal{I,F}\},\end{split}$ (4)
where $T_{\mathrm{c}}$ is the channel coherence time, $\vartheta_{u}$ is the QoS exponent for queuing delay, $\mathbb{E}[\cdot]$ denotes the expectation, and $r_{u}^{D,\xi}$ is the data rate in (2) or (3). We note that $\vartheta_{u}=\frac{\ln(1/\varepsilon_{u})}{a_{u}{\tau}_{\max}}$ is determined by the maximum tolerable delay bound violation probability, $\varepsilon_{u}$, the packet arrival rate, $a_{u}$, and the threshold of queuing delay, $\tau_{\max}$.
III-A3 Security Requirement
To formulate the wireless security requirement, we consider that there is an eavesdropper that attempts to wiretap the information transmitted by each user.
@@ III-A3 Security Requirement
where $[x]^{+}=\max\{0,x\}$, and $r_{u}^{e,\mathcal{I}}=\frac{w_{u}}{\ln 2}\ln\left(1+\frac{P_{u}h_{u}^{e}}{w_{u}N_{0}}\right)$ is the data rate of the wiretapped channel from the $u$-th user to the eavesdropper. The channel gain of the wiretapped channel is denoted by $h_{u}^{e}=\alpha_{u}^{e}g_{u}^{e}$, where ${\alpha}_{u}^{e}$ and $g_{u}^{e}$ represent the large-scale and small-scale channel gains between the $u$-th user and the eavesdropper, respectively.
@@
III-B Bandwidth Reservation for Different Slices
We assume that there can be multiple bandwidth reservation policies for different slices in network slicing. Given the total bandwidth of the BS, $W_{\max}$, the bandwidth reserved for the $\tau$-th slice is given by
where $U_{\tau}^{\Phi,\xi}$ is the number of users in the $\tau$-th slice, $I_{u_{\tau}}^{\Phi,\xi}$ is the QoS class identifier (QCI) of the $u_{\tau}$-th user in the $\tau$-th slice, and $f_{\tau}^{\mathrm{NS}}\left(\cdot,\cdot\right)$ is the network function for bandwidth reservation in network slicing. Since the sum of the bandwidth reserved for all the slices equals the total bandwidth of the BS, thus
where $T_{\mathrm{NS}}$ is the number of slices. Inspired by [31], the bandwidth reserved for each slice depends on the number of users in this slice and the QCI of these users, e.g.,
@@ III-C Problem Formulation
To maximize the sum reward subject to the QoS requirements in each slice, we formulate the bandwidth allocation problem as follows,
@@
where $\bm{w}_{\tau}^{\Phi,\xi}=[w_{1}^{\Phi,\xi},w_{2}^{\Phi,\xi},\cdots,w_{U_{\tau}}^{\Phi,\xi}]^{\mathrm{T}}$ is the bandwidth allocated to the users, and $r_{\tau}^{\Phi,\xi}$ is the minimum threshold of the QoS required by the users. Thus, constraint (10c) guarantees the QoS of all the users.
@@ III-D Analysis of Problem Feasibility
Given the available bandwidth constraint in (10a) and the QoS constraint in (10c), problem (10) will be infeasible when some of the users in this slice have weak channels. We denote the minimum bandwidth required to meet constraint (10c) by ${\bm{w}}_{\min}^{\Phi,\xi}=\left[{w}_{1,\min}^{\Phi,\xi},\cdots,{w}_{U_{\tau},\min}^{\Phi,\xi}\right]^{\mathrm{T}}$. If some users experience deep fading, leading to $\sum_{u\in\mathcal{U}_{\tau}^{\Phi,\xi}}{w}_{u,\min}^{\Phi,\xi}>W_{\tau}^{\Phi,\xi}$, then problem (10) is infeasible. In this case, the BS will only schedule the users with sufficiently strong channels. Alternatively, to maximize the number of scheduled users in problem (10), we consider that the BS schedules the $K$ users with the smallest bandwidth requirement. Denote the set of scheduled users by $\mathcal{K}_{\tau}^{\Phi,\xi}$. Then, for any $k\in\mathcal{K}_{\tau}^{\Phi,\xi}$ and $u\notin\mathcal{K}_{\tau}^{\Phi,\xi}$, we have ${w}_{k,\min}^{\Phi,\xi}\leq{w}_{u,\min}^{\Phi,\xi}$. After user scheduling, problem (10) can be reformulated as follows,

### v3

# Hybrid-Task Meta-Learning: A GNN Approach for Scalable and Transferable Bandwidth Allocation

@@ Abstract
In this paper, we develop a deep learning-based bandwidth allocation policy that is: 1) scalable with the number of users and 2) transferable to different communication scenarios, such as non-stationary wireless channels, different quality-of-service (QoS) requirements, and dynamically available resources. To support scalability, the bandwidth allocation policy is represented by a graph neural network (GNN), with which the number of training parameters does not change with the number of users. To enable the generalization of the GNN, we develop a hybrid-task meta-learning (HML) algorithm that trains the initial parameters of the GNN with different communication scenarios during meta-training. Next, during meta-testing, a few samples are used to fine-tune the GNN with unseen communication scenarios. Simulation results demonstrate that our HML approach can improve the initial performance by $8.79\%$, and sample efficiency by $73\%$, compared with existing benchmarks. After fine-tuning, our near-optimal GNN-based policy can achieve close to the same reward with much lower inference complexity compared to the optimal policy obtained using iterative optimization. Numerical results validate that our HML can reduce the computation time by approximately $200$ to $2000$ times than the optimal iterative algorithm.
@@ I Introduction
When the services with diverse QoS requirements co-exist in a network, the first step is to reserve physical resources among different types of services by network slicing (NS) technology for each slice [20], aiming to support diverse QoS requirements, such as data rate [22, 21], latency [23, 24], and security [26, 27, 25, 28], in both long and short coding blocklength regimes [31, 29, 30]. To reserve resources for a single slice, the authors of [32] proposed to compute the weights of different slices based on the corresponding QoS requirements and the number of service requests. With this approach, the amount of reserved resources for each slice is stochastic. Meanwhile, since the wireless channels are non-stationary, the reserved resources and the wireless channels in the training stage could be different from the actual required resources in the testing stage [33, 34]. As such, the mismatch between training data samples and testing data samples remains a crucial bottleneck for implementing efficient learning-based policies in practical wireless networks.
Recent works have considered to reduce the online training time using transfer learning, which involves offline pre-training of the neural network in a known scenario and online fine-tuning to apply it to a related scenario [10]. Compared with conventional random initialization, transfer learning can enhance sample efficiency in a target wireless communication scenario by leveraging well-trained neural network features from a related communication scenario. However, when considering a large number of communication scenarios including the inherent non-stationary nature of wireless communications, transfer learning will require more time to train the neural network due to limited data samples in each communication scenario. To further improve the online training efficiency for unseen wireless communication scenarios, meta-learning has been proposed [35, 38, 36, 37]. One of the meta-learning algorithms, model-agnostic meta-learning (MAML), has been applied to solve policy mismatch issues caused by varying user requests and non-stationary wireless channels [40, 39, 8, 9]. While these aforementioned works have highlighted the generalization ability of meta-learning for non-stationary wireless resource allocation, no works have addressed the impact of diverse QoS requirements in wireless communications.
@@
  • Finally, we develop our generalized hybrid-task meta-learning (HML) algorithm that is transferable to different communication scenarios by using meta-training to train the initial parameters of the GNN. We note that only a few samples are required to fine-tune the parameters of the GNN in meta-testing which validates that our GNN-based policy initialized by HML can be efficiently transferred to previously unseen communication scenarios. Simulation results show that our GNN-based policy achieves near-optimal performance and HML significantly outperforms the three considered benchmarks of MAML, MTL transfer (multi-task learning based transfer learning), and random initialization.
In our simulations, the sum reward achieved by the GNN-based policy and that of the optimal bandwidth allocation policy obtained from the iterative optimization algorithm can be overlapped in the long blocklength regime. The performance gap in between is less than $6\%$ even in the scenario considering effective capacity that the iterative algorithm has considerably higher computational complexity. HML also improves the initial performance by up to $8.79\%$ and sample efficiency by up to $73\%$ compared with the MAML benchmark. We also show that the performance gains of HML is even higher when compared to the other two benchmarks.
@@ II Related Works
II-A Deep Learning for Resource Allocation in Wireless Networks
Applying deep learning for resource allocation in wireless networks has been widely studied in the existing literature [15, 16]. In [15], the authors showed that learning-based algorithms could obtain near-optimal solutions by using a deep FNN, and the computational complexity in inference is low. In [16], the authors proposed an FNN-based unsupervised learning algorithm to optimize the bandwidth allocation policy. More recently, due to the fact that FNN is not scalable to the number of users, GNNs have been applied in wireless networks optimizations [19, 18]. In [18], the authors designed a GNN, which is scalable to the number of users in a wireless network, to minimize the summation of queuing delay violation probability and packet loss probability. In [19], the authors developed GNN-based scalable learning-based methods to solve radio resource management problems. How to efficiently generalize a scalable GNN to allocate bandwidth resources across different communication scenarios remains an open issue.
@@ II-A Deep Learning for Resource Allocation in Wireless Networks
Refs QoS
Data rate
Latency
Security
Long
Short
Long
Short
Long
Short
@@ II-B Generalization of Deep Learning Policies in Non-Stationary Wireless Networks
In wireless networks, the user requests, wireless channels, and available resources for each type of service can be non-stationary. Table I summarizes some QoS requirements considered in the related works. For example, data rate, latency, and security have been investigated in [21, 22, 23, 24, 25, 26, 27]. These papers mainly focus on scenarios with long channel coding blocklengths, where the achievable rate of a wireless link can be approximated by the Shannon capacity. In 5G communications, the coding blocklength can be short, and Shannon capacity is not applicable. As such, the authors of [28, 29] established to optimize wireless communication systems using the achievable rate in the short blocklength regime [30]. In addition, since different services may co-exist in one network, the authors of [31, 10] considered different QoS requirements in both long and short blocklength regimes. To support diverse QoS requirements in network slicing, the authors of [32] proposed to reserve bandwidth for different slices based on the number of users and the required QoS. As shown in Table I, none of the previously mentioned works proposed a general solution that can simultaneously support data rate, latency, and security in both long and short blocklength regimes.
Considering that the number of requests, the reserved resources, and the wireless channels all can be dynamic, improving the generalization ability of deep learning policies has attracted significant research interests in recent years. One approach to address this challenge is to deliberately initialize the neural network and fine-tune it online. The authors of [10] applied transfer learning to fine-tune the parameters of deep neural networks that are trained offline in dynamic wireless networks, thus to improve the sample efficiency when the communication scenario changed to the other one. To further improve the sample efficiency across different unseen communication scenarios, meta-learning has been adopted in [8, 9, 39, 40], where the hyper-parameters of a deep neural network, such as the initial parameters, are updated according to a set of communication scenarios in meta-training. In [39], meta-learning was applied to optimize computing resource allocation policies in mobile edge computing networks to fit both time-varying wireless channels and different requests of computing tasks. In [40], meta-learning was applied in virtual reality to quickly adapt to the user movement patterns changing over time. To improve the training efficiency in non-stationary vehicle networks, the authors in [8] proposed optimizing the beamforming using meta-learning. In [9], the authors combined meta-learning and support vector regression to extract the features for beamforming optimization, further improving training efficiency over non-stationary channels. There is no existing literature for using GNN-based meta-learning to implement network slices that support diverse communication scenarios.
@@ III System Model and Problem Formulation
III-A Bandwidth Reservation for Individual Slices
We assume that there can be multiple bandwidth reservation policies for different slices in network slicing (NS), and users of the same QoS requirement and the same blocklength share the same slice. We further use the QoS class identifier (QCI) to characterize the users in the slices [32]. Given the total bandwidth of the BS, $W_{\max}$, the bandwidth reserved for the $\tau$-th slice is given by
where superscript $\Phi$ represents the QoS requirement in the current network slice, whilst the superscript $\xi$ represents whether the current scenario is in a long or short blocklength regime, respectively. Function $f_{\tau}^{\mathrm{NS}}\left(\cdot,\cdot\right)$ is the NS function for bandwidth reservation in NS, $U_{\tau}^{\Phi,\xi}$ is the number of users in the $\tau$-th slice with QoS requirement as $\Phi$ and in long or short blocklength regime represented by $\xi$, and $I_{u_{\tau}}^{\Phi,\xi}$ is the QCI of the $u_{\tau}$-th user in the $\tau$-th slice. Since the sum of the bandwidth reserved for all the slices equals the total bandwidth of the BS, thus can be described as
where $T_{\mathrm{NS}}$ is the number of total network slices managed by the BS. Inspired by [32], the bandwidth reserved for each slice depends on the number of users in this slice and the QCI of these users, e.g.,
III-B Different QoS in Long and Short Blocklength Regimes
To investigate the generalization ability of our proposed bandwidth allocation policy, we consider both long and short blocklength regimes with three types of QoS requirements, i.e., data rate, queuing delay, and security. Thus, there are six communication objectives in total. We present details of these six objectives from the following three aspects.
III-B1 Data Rate Requirement
When the blocklength is long, the data rate reward of the $u$-th user can be expressed as [10]
@@ III-B1 Data Rate Requirement
III-B2 Latency Requirement
When considering latency constraints due to queueing delays, the effective capacity is applied to characterize the statistical QoS requirement in wireless communications. The effective capacities in long and short blocklength regimes are expressed as  [29]
$r_{u}^{E,\mathcal{I}}=-\frac{1}{\vartheta_{u}T_{\mathrm{c}}}\ln\left(\mathbb{E}_{g_{u}}\left[\exp\left(-\vartheta_{u}T_{\mathrm{c}}r_{u}^{D,\mathcal{I}}\right)\right]\right),$ (6)
and
$r_{u}^{E,\mathcal{F}}=-\frac{1}{\vartheta_{u}T_{\mathrm{c}}}\ln\left(\mathbb{E}_{g_{u}}\left[\exp\left(-\vartheta_{u}T_{\mathrm{c}}r_{u}^{D,\mathcal{F}}\right)\right]\right),$ (7)
respectively, where $T_{\mathrm{c}}$ is the channel coherence time, $\vartheta_{u}$ is the QoS exponent for queuing delay, $\mathbb{E}[\cdot]$ denotes the expectation, and $r_{u}^{D,\xi}$ is the data rate in (4) or (5). We note that $\vartheta_{u}=\frac{\ln(1/\varepsilon_{u})}{a_{u}{\tau}_{\max}}$ is determined by the maximum tolerable delay bound violation probability, $\varepsilon_{u}$, the packet arrival rate, $a_{u}$, and the threshold of queuing delay, $\tau_{\max}$.
III-B3 Security Requirement
To formulate the wireless security requirement, we consider that there is an eavesdropper that attempts to eavesdrop the information transmitted by each user.
@@ III-B3 Security Requirement
where $x]^{+}=\max\{0,x\}$, and $r_{u}^{e,\mathcal{I}}=\frac{w_{u}}{\ln 2}\ln\left(1+\frac{P_{u}h_{u}^{e}}{w_{u}N_{0}}\right)$ is the data rate of the eavesdropped channel from the $u$-th user to the eavesdropper. The channel gain of the eavesdropped channel is denoted by $h_{u}^{e}=\alpha_{u}^{e}g_{u}^{e}$, where ${\alpha}_{u}^{e}$ and $g_{u}^{e}$ represent the large-scale and small-scale channel gains between the $u$-th user and the eavesdropper, respectively.
@@ III-C Problem Formulation
To maximize the sum reward subject to the QoS requirements and blocklength in each slice, we formulate the bandwidth allocation problem as follows,
@@
where $\bm{w}_{\tau}^{\Phi,\xi}=[w_{1}^{\Phi,\xi},w_{2}^{\Phi,\xi},\cdots,w_{U_{\tau}}^{\Phi,\xi}]^{\mathrm{T}}$ is the bandwidth allocated to the users with QoS requirement and $r_{\tau}^{\Phi,\xi}$ is the minimum threshold of the QoS required by the users, where superscripts $\Phi$ represents the QoS requirement in the current network slice, and $D,E,S$ indicate data rate, effective capacity with queuing delay constraint, and secrecy rate, respectively; whilst the superscripts $\xi$ represents whether the current scenario is in a long or short blocklength regime, and $\mathcal{I,F}$ indicate the scenarios in the infinite long and finite short blocklength regimes, respectively. Thus, constraint (10c) guarantees the QoS of all the users. Taking eqs. (4)–(9) into problem (10a), we can observe that problem (10) is non-convex, since four out of the six communication scenarios are non-convex.
1 Initialize the set of scheduled users as all the users: $\mathcal{K}_{\tau}^{\Phi,\xi}=\mathcal{U}_{\tau}^{\Phi,\xi}$.
2 1) Drop $u$-th user if the target in eq. (10c) cannot be satisfied with $W_{\tau}^{\Phi,\xi}$:
3 for $u\in\mathcal{U}_{\tau}^{\Phi,\xi}$ do
4  Calculate the reward of each user by taking $w_{u}=W_{\tau}^{\Phi,\xi}$ into the target reward shown in eqs. (4)–(9) as: $r_{u,\max}^{\phi,\xi}=r_{u}^{\phi,\xi}(W_{\tau}^{\Phi,\xi})$.
5  if $r_{u,\max}^{\phi,\xi}<r_{\tau}^{\Phi,\xi}$ then
6     Drop the $u$-th user: $\mathcal{K}_{\tau}^{\Phi,\xi}=\mathcal{K}_{\tau}^{\Phi,\xi}-\{u\}$, K=K1K=K-1, and $w_{u}=0$.
7
8  else
9    Get $w_{k,\min}$ using binary search function: $w_{k,\min}=f_{\mathrm{BiS}}(h_{k},r_{\tau}^{\Phi,\xi})$.
10  end if
11
12 end for
132) Drop $k$-th user if eq. (10a) cannot be satisfied with $w_{k,\min}$:
14 while $\sum\nolimits_{k=1}^{K}{w_{k,\min}}>W_{\tau}^{\Phi,\xi}$  do
15  Identify user holding highest $w_{k,\min}$ in $\mathcal{K}_{\tau}^{\Phi,\xi}$: $k_{\mathrm{drop}}=\arg\max\limits_{k}{w_{k,\min}}$.
16  Drop the $k_{\mathrm{drop}}$-th user: $\mathcal{K}_{\tau}^{\Phi,\xi}=\mathcal{K}_{\tau}^{\Phi,\xi}-\{k_{\mathrm{drop}}\}$, K=K1K=K-1, and $w_{k_{\mathrm{drop}}}=0$
17
18 end while
Algorithm 1 User Scheduling Algorithm
@@ III-D Analysis of Problem Feasibility
Given the available bandwidth constraint in (10a) and the QoS constraint in (10c), problem (10) will be infeasible when some of the users in this slice have weak channels. We denote the minimum bandwidth required to meet constraint (10c) by ${\bm{w}}_{\min}^{\Phi,\xi}=\left[{w}_{1,\min}^{\Phi,\xi},\cdots,{w}_{U_{\tau},\min}^{\Phi,\xi}\right]^{\mathrm{T}}$. To get the minimum required bandwidth, we first assess if each user’s reward meets constraint (10c) by line 1 of Algorithm 1. If not, it indicates this user is experiencing a deep fading channel, necessitating infinite bandwidth. In contrast, for users not experiencing deep fading, we employ a binary search function to compute their required minimum bandwidth. The step-by-step process is detailed in lines 1–10 of Algorithm 1. However, it is possible that the sum minimum bandwidth achieved in line 8 of Algorithm 1 is larger than $W_{\tau}^{\Phi,\xi}$, which is conflicted with constraint (10a) conflicting with constraint (10a). Therefore, we must remove the users sequentially consuming the most bandwidth until constraint (10a) is satisfied. We note that if some users experience deep fading, leading to $\sum_{u\in\mathcal{U}_{\tau}^{\Phi,\xi}}{w}_{u,\min}^{\Phi,\xi}>W_{\tau}^{\Phi,\xi}$, then problem (10) is infeasible. In this case, the BS will only schedule the users with sufficiently strong channels. Alternatively, to maximize the number of scheduled users in problem (10), we consider that the BS schedules the $K$ users with the minimum bandwidth requirement. Denote the set of scheduled users by $\mathcal{K}_{\tau}^{\Phi,\xi}$. Then, for any $k\in\mathcal{K}_{\tau}^{\Phi,\xi}$ and $u\notin\mathcal{K}_{\tau}^{\Phi,\xi}$, we have ${w}_{k,\min}^{\Phi,\xi}\leq{w}_{u,\min}^{\Phi,\xi}$. To achieve the scheduled users, we need to execute a user scheduling algorithm, and the step-by-step process is given in Algorithm 1. With Algorithm 1, we can optimize the number of scheduled users with the changes in the total number of requesting users. However, it’s also practical to consider alternative user scheduling algorithms for different requirements, such as prioritizing users based on their historical usage patterns, implementing a round-robin scheduling approach to ensure equal distribution of resources among users, or incorporating machine learning techniques to adjust schedulings based on real-time traffic and user behavior dynamically.
After user scheduling, problem (10) can be reformulated as follows,